I am having so much trouble trying to remember what makes an integer and what makes an irrational number, is there some cute little thing similar to PEMDAS that might help me?
Answer by jim_thompson5910(35256)

Integer: Any whole number like 2, 7, 11, etc


Rational number: any number that can be represented by a fraction of integers (eg: 2/3, 4/5, 0.4 since 0.4 = 4/10 = 2/5)


Irrational number: any number that CANNOT be represented by a fraction of integers (ex: pi = 3.14159265... it goes on forever and you cannot represent pi with a fraction of integers, ex: sqrt(2) = 1.4142135623731... same story with this one as well, you cannot represent sqrt(2) with a fraction of integers)